OpenStudy (anonymous):

choose x to max log(1-x)

OpenStudy (anonymous):

k so u cant take log of a negative number .... its a property will be helpful to memorize it .... max x = 1

OpenStudy (anonymous):

right..how would i solve with first order conditions?

OpenStudy (anonymous):

it has no maximum value.. it goes to infinity as x goes to negative infinity

OpenStudy (anonymous):

can you solve this with derivatives?

OpenStudy (anonymous):

yeah the derivative of log(1-x)=-1/(1-x) when you set it equal to zero, it has no solution.. therefore, there are no extreme values to the given function

OpenStudy (anonymous):

i see that makes sense, however in a similar problem the answer says x = 0, i.e., choose x to maximize (y - x)

OpenStudy (anonymous):

sorry the question is choose x to maximize log(y-x)

OpenStudy (anonymous):

are you sure about the answer? that it's x=0

OpenStudy (anonymous):

because I think negative values of x give greater values for the function log(y-x) than x=0

OpenStudy (anonymous):

oh i think they are assuming that x greater than or equal to zero

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

then i'm not sure how the derivate would yield x=0

OpenStudy (anonymous):

any ideas?